OpenStudy (anonymous):

NEED HELP ASAP! What is the slope of the line passing through the points (0, 4) and (–8, –1)? A. B. C. D.

OpenStudy (anonymous):

A -8/5 B -5/8 C 5/8 D 8/5

OpenStudy (emmamink):

Ok. slope is\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1}}\]

OpenStudy (anonymous):

ok

OpenStudy (emmamink):

so -1 - 4 is -5 and -8 minus 0 is -8. So we have -5/-8 as our answer.

OpenStudy (emmamink):

or just -5/8

OpenStudy (anonymous):

ok thanks can u please help me with one more

OpenStudy (emmamink):

Medal if you want. I also need fans

OpenStudy (emmamink):

sure

OpenStudy (emmamink):

post away and tag me

OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (emmamink):

Ok, look at the formula above.

OpenStudy (emmamink):

Let's plug in the values

OpenStudy (emmamink):

7 minus 3 is 4 and 4 minus 1 is 3 so we have 4/3

OpenStudy (anonymous):

a 2 b 3/4 c -4/3 d -2

OpenStudy (anonymous):

on second i think i rote it rong

OpenStudy (emmamink):

C. sorry i forgot to add the nagative sign

OpenStudy (anonymous):

What is the slope of the line passing through the points -1,3 4,-7

OpenStudy (emmamink):

so ya it's C.

OpenStudy (anonymous):

thank so much can u help with one more?

OpenStudy (anonymous):

What is the slope of the line whose equation is 4x – 6y = 12?

OpenStudy (anonymous):

4x-6y=12

OpenStudy (emmamink):

ok u know how to do it. I'm not doing all your test problems

OpenStudy (anonymous):

i think its 2/3

OpenStudy (emmamink):

Figure out 2 values for x and y, then do the formula

OpenStudy (emmamink):

it's pretty simple if u think about it.

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

is it rite?

OpenStudy (emmamink):

hold on....

OpenStudy (emmamink):

yes it is correct! Good job!

OpenStudy (anonymous):

thanks!! :) :)

OpenStudy (anonymous):

bye :) btw i love the cat

OpenStudy (emmamink):

thanks! can u fan me

OpenStudy (anonymous):

i did

OpenStudy (emmamink):

oh thanks!

OpenStudy (emmamink):

bye!